17. What is T, and T2 if the pulleys are frictionless and the coefficient of kinetic friction between the block and the tableIf you can explain each step I would appreciate it! Thanks.
Answer #1

FBD for 3kg block on table: 3a 30 a A TI MENT 39 Balancing forces in x direction: T2 - T, - HAN - 39 sin 30 = 3a Balancing foFBD for 3kg hanging block! | FBD for 8 kg hanging block: 3a 131 39 T, -39 = 3a T, = 30 + 3a 8g - T2 = 8a T2 = 89 80-8a [from